OpenStudy (anonymous):

Plant A is 18 inches tall after 1 week, 36 inches tall after two weeks, 56 inches tall after three weeks. Plant B is 18 inches tall after one week, 36 inches tall after two weeks, 54 inches tall after three weeks. which situation represents a proportional relationship between the plants' height and number of weeks? (How can you show the work in a graph?)

OpenStudy (anonymous):

@satellite73 Do you know this?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

Ok thanks anyway @satellite73

OpenStudy (texaschic101):

Not too sure about a graph... proportions are nothing but equivalent fractions. 1/18 = 2/36 ...but does not equal 3/56....so the first one is not proportional 1/18 = 2/36 = 3/54.....these all reduce to 1/18...so this is proportional

OpenStudy (anonymous):

Thank You! :D

OpenStudy (texaschic101):

no problem...sorry about the graph thing :)

OpenStudy (anonymous):

Its ok thnx tho :D